Operations Report – April 07th

Crew 191 Operations Report 07/04/2018

SOL: 14

Name of person filing report: Wataru Okamoto

Non-nominal systems: Water Tank

Notes on non-nominal systems: We cleaned up all tank. Loft water filter is already ordered by Dr. Shannon. Now water line is not filter line,another line bypassed.

Generator (hours run): 12hours

Solar SOC 83% (Before generator is run at night)

Diesel 80%

Propane 70psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) ? gallons.

Water (trailer) 0%

Water (static) 60%

Trailer to Static Pump used ? Yes

Water (loft) Static to Loft Pump used Yes

Water Meter: 133208.1
Toilet tank emptied: No

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) –

Oil Added –

ATV Fuel Used:
Hours the ATVs were Used today:

Notes on ATVs:

Deimos rover used: Yes
Hours: 118.6
Beginning charge: %
Ending charge: %
Currently charging: plugged

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: Yes
Hours: 30.2h
Beginning charge: 100%
Ending charge: 100%
Currently charging:

Opportunity rover used: –
Hours: 28.8h
Beginning charge: 100%
Ending charge: 100%
Currently charging:

Curiosity rover used: Yes

Hours: 25.1h
Beginning charge: 100%
Ending charge: 100%
Currently charging: plugged

HabCar used and why, where? We went to Hanksville 3 times. Because water supply.

General notes and comments: We explain to CREW192 about engineering.

Summary of internet: No problem

Summary of suits and radios: 5 charger of the backpack out of order.

Summary of Hab operations: We supply the black tank 20%.

Summary of GreenHab operations: No problem

Summary of ScienceDome operations: CREW192 member will sleep inside the dome tonight.

Summary of RAMM operations: N/A

Summary of health and safety issues: No Problem

Questions, concerns and requests to Mission Support: Thank you for your support. We completed the mission.